OpenStudy (anonymous):

Please help!! log8(x+2)-log8(x-1)=2/3 Since the base is common I set log8 (x+2)(x-1)=2/3. CAn anyone let me know if this is correct and help me find the solution, please. Thanks in advanced.

OpenStudy (asnaseer):

almost - the correct rule to use here would be:\[\log_ap-\log_aq=\log_a(\frac{p}{q})\]

OpenStudy (asnaseer):

you can find some basic log rules here to help you: http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

so that would be log8 x+2/x-1=2/3??

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

i have the rules....i'm just getting stumped on the steps. I need a little guidance.

OpenStudy (asnaseer):

ok - np

OpenStudy (asnaseer):

so do you know what the next step should be?

OpenStudy (asnaseer):

remember, if:\[\log_ap=n\implies p=a^n\]

OpenStudy (anonymous):

2/3(x-1)=(x+2)??

OpenStudy (asnaseer):

no - look carefully at the rule I just posted above and try it again

OpenStudy (asnaseer):

in your case:\[a=8\]\[p=\frac{x+2}{x-1}\]\[n=\frac{2}{3}\]

OpenStudy (anonymous):

ok.....working it out now...hold on a sec. thanks

OpenStudy (asnaseer):

np - take your time :)

OpenStudy (anonymous):

i got log8 x+2/x-1=2/3

OpenStudy (anonymous):

i now multiply 2/3 by both sides to get rid of teh fraction correct? or is it the inverse 3/2

OpenStudy (asnaseer):

no, after your first simplification to:\[\log_8(\frac{x+2}{x-1})=\frac{2}{3}\]you need to use the second rule I gave you, i.e. if:\[\log_ap=n\]then this implies:\[p=a^n\]

OpenStudy (anonymous):

(x+2/x-1)=8^2/3

OpenStudy (asnaseer):

correct

OpenStudy (asnaseer):

now try and simplify the right-hand-side, i.e. simplify:\[8^{\frac{2}{3}}\]

OpenStudy (asnaseer):

this can be written as:\[8^{\frac{2}{3}}=(8^{\frac{1}{3}})^2\]

OpenStudy (asnaseer):

and \(8^{\frac{1}{3}}\) means the cube-root of 8

OpenStudy (anonymous):

not (2^3)^2/3

OpenStudy (asnaseer):

where did you get that from?

OpenStudy (anonymous):

lol....my brain is so mixed up...sorry

OpenStudy (asnaseer):

:)

OpenStudy (asnaseer):

the last correct step we got to was:\[\frac{x+2}{x-1}=8^{\frac{2}{3}}\]so then I asked you to first simplify the right-hand-side of this equation, i.e. simplify:\[8^{\frac{2}{3}}=(8^{\frac{1}{3}})^2\]

OpenStudy (asnaseer):

where \(8^{\frac{1}{3}}\) means the cube-root of 8

OpenStudy (asnaseer):

do you know what the cube-root of 8 is?

OpenStudy (anonymous):

2

OpenStudy (asnaseer):

good, so we can write:\[8^{\frac{2}{3}}=(8^{\frac{1}{3}})^2=(2)^2=?\]

OpenStudy (anonymous):

4

OpenStudy (asnaseer):

great, so now going back to our last equation we get:\[\frac{x+2}{x-1}=8^{\frac{2}{3}}=4\]next step is to multiply both sides by (x-1)

OpenStudy (anonymous):

answer is x=2

OpenStudy (asnaseer):

correct

OpenStudy (anonymous):

yes...site froze

OpenStudy (anonymous):

i worked it out...i wish i had an answer sheet

OpenStudy (asnaseer):

ok - thats even better! these types of problems just need plenty of practice - you'll be a pro pretty soon!

OpenStudy (anonymous):

thank you soo much.i hope so

OpenStudy (asnaseer):

yw